# Are Economists Always More Introverted? Analyzing Consistency in Persona-Assigned LLMs

### v1

@@ Abstract
Manon Reusens1, Bart Baesens1,2 and David Jurgens3 1Research Centre for Information Systems Engineering (LIRIS), KU Leuven 2Department of Decision Analytics and Risk, University of Southampton 3 School of Information, University of Michigan manon.reusens@kuleuven.be, bart.baesens@kuleuven.be, jurgens@umich.edu
@@ 1 Introduction
We focus on two primary aspects of consistency: (1) Intra-persona consistency — whether LLMs remain consistent within their assigned persona; and (2) Inter-persona consistency — whether LLMs remain consistent across other persona categories than the one assigned.
@@ 2.1 Persona construction
The left part of the consistency framework shown in Figure 1 illustrates the persona construction. We selected the persona categories based on three key criteria: relevance in persona-related literature, variability in the scale of linguistic expression, and the availability of external survey instruments. Specifically, we focused on personality, professions, and political stance Malik et al. (2024); Jiang et al. (2024); Wang et al. (2024b), which are well-studied categories in the persona-related literature. Additionally, we included a binary persona category—happy or sad—as a useful contrast, since emotional states tend to be more explicitly reflected in language, whereas categories like occupation may manifest more subtly. This range of personas allows us to examine varying degrees of consistency. Finally, for each identified persona category, we defined personas based on well-known surveys:
@@
Based on the outcomes of the surveys, we constructed the different personas by making all possible combinations of the persona characteristics within a persona category, e.g. for the political category we include economically left and socially libertarian; economically right and socially libertarian; economically left and socially authoritarian; economically right and socially authoritarian. Additional information on the selection criteria for the persona categories are provided in Appendix A and all personas are included in Appendix B. This comprehensive approach ensures that our framework captures a wide range of realistic and nuanced persona scenarios.
@@ 2.3 Scoring Key
Additional information on the scoring keys is provided below. A more detailed explanation, including the prompts and specific survey scoring mechanisms, can be found in Appendix C.
@@
For the other dimensions, we use an LLM-as-a-judge, GPT4o, to evaluate the final outcomes, determining the persona’s alignment with binary characteristics, or for occupations one of the six different classes. The evaluation process is consistent across all dimensions, with the LLM assessing all characteristics across all responses. Detailed information on the used prompt is provided in Appendix C.2. The model also provides a confidence score on a four-point Likert scale per choice. A neutral choice was identified when the model’s confidence score was 1 or 2. To validate the reliability of the LLM judgments, we manually annotated 100 examples across all evaluation categories and found a Cohen’s $\kappa$ of 0.68 with these final LLM results supporting the reliability of the LLM-generated judgments.
@@ 2.4 Consistency scores
To measure consistency, we use Shannon entropy (Shannon, 1948), a metric that captures the uncertainty in the distribution of predicted labels across responses. It applies to both binary and multiclass persona traits and reflects full distributional patterns rather than just majority labels. Crucially, entropy allows us to compare consistency across models and persona categories without relying on arbitrary thresholds.
@@ 3.2 Experimental set-up
We analyze the consistency over 5 runs across 5 models from 3 different model families: Qwen-2.5 32B, Ministral-8B, Llama-3.2 3B, Llama-3.1 8B, and Llama-3.3 70B. Additional information on the checkpoints used is included in Appendix D We analyze the entropy per model and per evaluation and persona category. To gather insights into the chosen labels per characteristic, we examine the characteristic-specific consistency. Next, we analyze the overall consistency making cross-dimension comparisons.
@@ 3.3 Results
Happiness $0.18\pm 0.25$ $0.26\pm 0.41$ $0.14\pm 0.08$ $0.21\pm 0.44$
Occupation $0.59\pm 0.39$ $0.21\pm 0.21$ $0.52\pm 0.33$ $0.51\pm 0.31$
Personality $0.20\pm 0.14$ $0.15\pm 0.11$ $0.25\pm 0.15$ $0.22\pm 0.11$
Political $0.80\pm 0.45$ $0.76\pm 0.43$ $0.75\pm 0.40$ $0.41\pm 0.46$
@@
Happiness $0.26\pm 0.25$ $0.27\pm 0.38$ $0.38\pm 0.16$ $0.45\pm 0.36$
Occupation $0.76\pm 0.15$ $0.38\pm 0.31$ $0.71\pm 0.15$ $0.55\pm 0.28$
Personality $0.42\pm 0.15$ $0.24\pm 0.15$ $0.38\pm 0.11$ $0.34\pm 0.08$
Political $0.86\pm 0.19$ $0.92\pm 0.11$ $0.86\pm 0.16$ $0.51\pm 0.35$
@@
Happiness $00.00\pm 0$ 0.00 $0.30\pm 0.25$ $0.54\pm 0.12$ $0.64\pm 0.17$
Occupation $0.73\pm 0.16$ $0.42\pm 0.24$ $0.66\pm 0.21$ $0.63\pm 0.23$
Personality $0.31\pm 0.19$ $0.31\pm 0.12$ $0.42\pm 0.11$ $0.43\pm 0.16$
Political $0.84\pm 0.23$ $0.81\pm 0.25$ $0.89\pm 0.13$ $0.70\pm 0.18$
@@
Happiness $0.07\pm 0.16$ $0.21\pm 0.14$ $0.43\pm 0.09$ $0.39\pm 0.14$
Occupation $0.56\pm 0.30$ $0.23\pm 0.19$ $0.66\pm 0.24$ $0.59\pm 0.21$
Personality $0.30\pm 0.15$ $0.22\pm 0.04$ $0.35\pm 0.13$ $0.41\pm 0.12$
Political $0.83\pm 0.33$ $0.75\pm 0.38$ $0.79\pm 0.30$ $0.36\pm 0.31$
@@
Happiness $0.07\pm 0.15$ $0.05\pm 0.09$ $0.30\pm 0.19$ $0.15\pm 0.10$
Occupation $0.62\pm 0.38$ $0.23\pm 0.20$ $0.56\pm 0.34$ $0.49\pm 0.34$
Personality $0.23\pm 0.16$ $0.16\pm 0.09$ $0.27\pm 0.18$ $0.26\pm 0.13$
Political $0.79\pm 0.44$ $0.74\pm 0.43$ $0.71\pm 0.41$ $0.40\pm 0.31$
@@
Table 1: The tables show large entropy differences between the models, indicating differences in consistency levels. Both strong within-category consistency (diagonal) and occasional spill-over consistency (off-diagonal) are found, where lower is more consistent. Scores $<$0.25 are colored green, 0.25$<$0.5 in orange, and $>$0.5 in red. The columns represent the different assigned persona categories. The rows represent the evaluation categories. The standard deviation is computed over the entropy scores over different dimensions per evaluation-persona category pair.
@@
Figure 2: Qwen-2.5 32B generally follows the instructions and shows spill-over effects, i.e. stereotypes and default values. Columns denote personas, and rows indicate evaluation categories. Multi-component personas (e.g., political stance) are grouped per component and averaged over all personas containing that component. The other models are shown in Appendix E.
@@
As shown in Table 1, the diagonal values indicate relatively high consistency, meaning that when a persona is assigned, the model tends to generate responses that consistently express that specific persona across different output formats and prompts. However, the degree of consistency varies across persona categories. While happiness and occupation personas are more consistently expressed, personality and political personas exhibit lower intra-persona consistency. For the political category, we observe a high standard deviation, indicating substantial variability in consistency across dimensions. This is largely due to certain tasks, such as singlechat, where expressing a consistent political stance is more challenging in general. Manual analysis also confirms this finding, highlighting the difficulty of a model to express the political opinion when being asked certain questions. Here, the inconsistency thus stems from a lack of expression of the underlying persona. Similarly, personality-based personas show lower intra-persona consistency, which we investigate further in Figure 2. This figure shows the results for Qwen-2.5 32B (other models are shown in Appendix E). We chose Qwen as it provides representative results for all models with an average correlation of 0.70 with the other models. Examining Figure 2, we find that the LLM generally follows our instructions, e.g., the happy persona is more happy (1), while the sad persona is more sad (0). Likewise assigned occupations are clearly reflected in the output. However, certain personality traits—such as low conscientiousness, antagonism, neuroticism, and, in some models, low openness to experience—are more difficult for the LLM to express consistently, resulting in greater variability in responses within that personality category. Similarly, Salecha et al. (2024) show that models skew responses to socially desirable answers for the Big Five Personality test when they infer that they are evaluated. We demonstrate that social desirability bias also appears in other evaluation dimensions. This social desirability tendency explains the model’s difficulty to adhere to the unconscientious and antagonistic personas.
@@
The off-diagonal elements in the subtables of Table 1 reveal that most spill-over consistency effects occur across two evaluation categories: happiness and personality. For example, the off-diagonal entropy scores for Llama-3.3 70B are lower for the happiness and personality categories compared to occupation and political stance. This suggests that the model is more consistent across these two categories when they are not the assigned persona. Interestingly, this pattern differs from the intra-persona results, where happiness and occupation showed the strongest consistency when they were the assigned persona. The occupation and political categories are harder for the models to express, as not adding those personas results in responses without any occupational information or political stance. Manual analysis reveals how these personas are less frequently and explicitly expressed, especially in conversational settings. For instance, political beliefs rarely surface in responses to questions like "What are your music preferences?". Similarly, occupation-related information rarely appears unless explicitly prompted, though it may occasionally surface in essay-style or social media posts. The other two categories show spill-over effects. Additional manual analysis reveals that happiness and personality directly influence linguistic style—models default to a positive tone unless instructed otherwise, making happiness more overt. Similarly, personality traits, like extroversion, shape response style, amplifying spillover effects.
@@
Figure 2 shows how a sad persona is portrayed as more introverted, less conscientious, and less open than a happy persona. The sad persona is more likely to have an artistic occupation, while the happy persona is most likely to have a social occupation. The economically right-winged and socially authoritarian personas are both less agreeable than their counterparts. All occupation personas are presented as extroverts, except the economist, who is more introverted. These observations reinforce prior findings that persona-assigned LLMs are susceptible to stereotypes Gupta et al. (2023). We show that these stereotypes also appear across general text-generation tasks. Furthermore, the model tends to answer in a happy, conscientious, agreeable, and open manner unless otherwise instructed or influenced by a stereotype. This is reflected in the heatmaps: the rows corresponding to the happiness, conscientiousness, and agreeableness evaluation categories generally lean toward a value of 1, indicating strong alignment. In contrast, neuroticism tends to lean toward 0, suggesting lower identification with that trait. This again illustrates the social desirability bias in the models  Salecha et al. (2024). Additionally, we show how personas can partially counteract this bias, e.g. the high consistency scores for neurotic and sad personas. However, this effect is not universal, as evidenced by the lower intra-persona consistency for antagonistic and unconscientious personas. Furthermore, most models tend to be slightly economically left and socially libertarian, though this varies by model and sometimes leans toward inconsistency. These default personas reveal the LLM’s ideological stances, shaped by training data and choices from model developers, called design choices Buyl et al. (2025); Cambo and Gergle (2022).
@@
Figure 3 shows that the survey, essay, and social media dimensions have the lowest intra- and inter-persona entropy, indicating the highest consistency across these tasks. However, consistency scores vary notably across dimensions, with mostly low correlations between them. This highlights the importance of considering all three dimensions to fully capture model behavior. Of these, only the inter-persona consistency between essay and social media is strongly correlated. Their low entropy scores could be attributed to the clarity of the task, where models can easily express their assigned persona. As tasks become less straightforward—such as answering open-ended questions about music preferences (singlechat and multichat)—models generally show a decrease in both intra- and inter-persona consistency. Moreover, the difference between intra- and inter-persona entropy becomes smaller. These results suggest that as tasks require less explicit persona expression, models may struggle to express distinct persona characteristics. Depending on the application, this variability may be advantageous, i.e. in creative or open-ended generation tasks where diversity is desirable. However, in more controlled settings that require reliable persona adherence, such unpredictability can be a drawback, as the model’s outputs may deviate from the intended persona or produce inconsistent behavior. Interestingly, the complexity of multichat scenarios compared to singlechat does not appear to hinder consistency. Contrarily, consistency increases slightly as follow-up responses allow models to provide more information, expressing the assigned persona more clearly. Nevertheless, consistency scores between singlechat and multichat are highly correlated. In addition to the main results, supplementary statistical analyses revealed no significant difference between inter- and intra-persona consistency in the singlechat and multichat dimensions. However, we observed statistically significant differences between the evaluation dimensions of survey, essay, and social media on the one hand, and singlechat (for both intra- and inter-persona consistency) as well as multichat (for intra-persona consistency) on the other. Concerning inter-persona consistency specifically, only the social media post generation task significantly outperforms the inter-persona consistency observed in the multichat evaluation dimension. Detailed statistical analysis is reported in Appendix G.

### v2

@@ Abstract
Manon Reusens1,2, Bart Baesens1,3 and David Jurgens4 1Research Centre for Information Systems Engineering (LIRIS), KU Leuven 2Department of Engineering Management, University of Antwerp 3Department of Decision Analytics and Risk, University of Southampton 4 School of Information, University of Michigan
@@ 1 Introduction
We focus on two primary aspects of consistency: (1) Intra-persona consistency: whether LLMs remain consistent within their assigned persona; and (2) Inter-persona consistency: whether LLMs remain consistent across other persona categories than the one assigned.
@@ 2.1 Persona construction
The left part of the consistency framework shown in Figure 1 illustrates the persona construction. We selected the persona categories based on three key criteria: relevance in persona-related literature, variability in the scale of linguistic expression, and the availability of external survey instruments. Specifically, we focused on personality, professions, and political stance Malik et al. (2024); Jiang et al. (2024); Wang et al. (2024b), which are well-studied categories in the persona-related literature. Additionally, we included a binary persona category (happy or sad) as a useful contrast, since emotional states tend to be more explicitly reflected in language, whereas categories like occupation may manifest more subtly. This range of personas allows us to examine varying degrees of consistency. Finally, for each identified persona category, we defined personas based on well-known surveys:
@@
Additional information regarding the surveys used can be found in Appendix A. Based on the outcomes of the surveys, we constructed the different personas by making all possible combinations of the persona characteristics within a persona category, e.g. for the political category we include economically left and socially libertarian; economically right and socially libertarian; economically left and socially authoritarian; economically right and socially authoritarian. Additional information on the selection criteria for the persona categories are provided in Appendix B and all personas are included in Appendix C. This comprehensive approach ensures that our framework captures a wide range of realistic and nuanced persona scenarios.
@@ 2.3 Scoring Key
Additional information on the scoring keys is provided below. A more detailed explanation, including the prompts and specific survey scoring mechanisms, can be found in Appendix D.
@@
For the other dimensions, we use an LLM-as-a-judge, GPT4o, to evaluate the final outcomes, determining the persona’s alignment with binary characteristics, or for occupations one of the six different classes. The evaluation process is consistent across all dimensions, with the LLM assessing all characteristics across all responses. Detailed information on the used prompt is provided in Appendix D.2. The model also provides a confidence score on a four-point Likert scale per choice. A neutral choice was identified when the model’s confidence score was 1 or 2. To validate the reliability of the LLM judgments, we manually annotated 100 random examples across all evaluation categories from our dataset, with 25 samples per assigned persona category. This resulted in 50 sentences from singlechat, 42 from multichat, 4 from essay, and 4 from socialmedia. We acknowledge the distribution skews toward singlechat and multichat, which are notoriously challenging for consistency evaluations due to conversational variability and nuanced phrasing. We found a Cohen’s $\kappa$ of 0.68 with these final LLM results supporting the reliability of the LLM-generated judgments. Following  Landis and Koch (1977), this score lies within the substantial agreement range. Given that most of the validated samples represent the hardest categories for evaluation, this level of agreement is, in our opinion, a reasonably strong indication that the LLM’s judgments align with human intuitions under challenging conditions. Additionally, this task is already rather difficult for people. For example, Preoţiuc-Pietro et al. (2015) show how inferring subjective characteristics about users from socialmedia posts such as optimism, leads to cohen’s kappa ranges between 0.3 and 0.7.
@@ 2.4 Consistency scores
To measure consistency, we use Shannon entropy (Shannon, 1948), a metric that captures the uncertainty in the distribution of predicted labels across responses. It applies to both binary and multiclass persona traits and reflects full distributional patterns rather than just majority labels. Crucially, entropy allows us to compare consistency across models and persona categories without relying on arbitrary thresholds. Additional information about the choice of the metric is included in Appendix E
@@ 3.2 Experimental set-up
We analyze the consistency over 5 runs across 5 models from 3 different model families: Qwen-2.5 32B, Ministral-8B, Llama-3.2 3B, Llama-3.1 8B, and Llama-3.3 70B. Additional information on the checkpoints used is included in Appendix F We analyze the entropy per model and per evaluation and persona category. To gather insights into the chosen labels per characteristic, we examine the characteristic-specific consistency. Next, we analyze the overall consistency making cross-dimension comparisons.
@@ 3.3 Results
Happiness $0.28\pm 0.26$ $0.25\pm 0.41$ $0.15\pm 0.09$ $0.20\pm 0.44$
Occupation $0.55\pm 0.38$ $0.21\pm 0.21$ $0.50\pm 0.33$ $0.50\pm 0.33$
Personality $0.19\pm 0.13$ $0.16\pm 0.11$ $5\pm 0.15 0.22±0$$2\pm 0.11$
Political $0.80\pm 0.45$ $0.77\pm 0.43$ $0.75\pm 0.40$ $0.39\pm 0.46$
@@
Happiness $0.26\pm 0.25$ $0.28\pm 0.38$ $0.38\pm 0.17$ $0.40\pm 0.35$
Occupation $0.76\pm 0.18$ $0.40\pm 0.32$ $0.68\pm 0.16$ $0.57\pm 0.21$
Personality $0.42\pm 0.15$ $0.24\pm 0.15$ $0.39\pm 0.11$ $0.37\pm 0.08$
Political $0.91\pm 0.17$ $0.93\pm 0.07$ $0.85\pm 0.15$ $0.51\pm 0.35$
@@
Happiness $0.00\pm 0.00$ $0.26\pm 0.28$ $0.52\pm 0.11$ $0.65\pm 0.17$
Occupation $0.70\pm 0.13$ $0.38\pm 0.26$ $0.64\pm 0.21$ $0.62\pm 0.20$
Personality $\pm 0.18 0.31±0.13$$\pm 0.13$ $0.43\pm 0.11$ $0.47\pm 0.17$
Political $0.86\pm 0.22$ $0.81\pm 0.25$ $0.89\pm 0.13$ $0.75\pm 0.13$
@@
Happiness $0.07\pm 0.16$ $0.22\pm 0.14$ $0.42\pm 0.12$ $0.43\pm 0.10$
Occupation $0.67\pm 0.29$ $0.23\pm 0.17$ $0.63\pm 0.24$ $0.57\pm 0.21$
Personality $0.31\pm 0.15$ $0.24\pm 0.07$ $0.36\pm 0.13$ $0.41\pm 0.10$
Political $0.83\pm 0.33$ $0.76\pm 0.38$ $0.79\pm 0.30$ $0.35\pm 0.30$
@@
Happiness $07\pm 0.15 0.0$$08\pm 0.09$ $0.30\pm 0.18$ $0.20\pm 0.16$
Occupation $0.51\pm 0.38$ $0.22\pm 0.18$ $0.55\pm 0.33$ $0.46\pm 0.35$
Personality $0.23\pm 0.14$ $0.17\pm 0.10$ $0.27\pm 0.18$ $0.23\pm 0.14$
Political $0.79\pm 0.44$ $0.76\pm 0.43$ $0.71\pm 0.40$ $0.36\pm 0.30$
@@
Table 1: The tables show large entropy differences between the models, indicating differences in consistency levels. Both strong within-category consistency (diagonal) and occasional spill-over consistency (off-diagonal) are found, where lower is more consistent. Scores <<0.25 are colored green, 0.25<<0.5 in orange, and >>0.5 in red. The columns represent the different assigned persona categories. The rows represent the evaluation categories. The standard deviation is computed over the entropy scores over different dimensions per evaluation-persona category pair.
@@
Figure 2: Qwen-2.5 32B generally follows the instructions and shows spill-over effects, i.e. stereotypes and default values. Columns denote personas, and rows indicate evaluation categories. Multi-component personas (e.g., political stance) are grouped per component and averaged over all personas containing that component. The other models are shown in Appendix G.
@@
As shown in Table 1, the diagonal values indicate relatively high consistency, meaning that when a persona is assigned, the model tends to generate responses that consistently express that specific persona across different output formats and prompts. However, the degree of consistency varies across persona categories. While happiness and occupation personas are more consistently expressed, personality and political personas exhibit lower intra-persona consistency. For the political category, we observe a high standard deviation, indicating substantial variability in consistency across dimensions. This is largely due to certain tasks, such as singlechat, where expressing a consistent political stance is more challenging in general. Manual analysis also confirms this finding, highlighting the difficulty of a model to express the political opinion when being asked certain questions. Here, the inconsistency thus stems from a lack of expression of the underlying persona. Similarly, personality-based personas show lower intra-persona consistency, which we investigate further in Figure 2. This figure shows the results for Qwen-2.5 32B (other models are shown in Appendix G). We chose Qwen as it provides representative results for all models with an average correlation of 0.66 with the other models. Examining Figure 2, we find that the LLM generally follows our instructions, e.g., the happy persona is more happy (1), while the sad persona is more sad (0). Likewise assigned occupations are clearly reflected in the output. However, certain personality traits, such as low conscientiousness, antagonism, neuroticism, and, in some models, low openness to experience, are more difficult for the LLM to express consistently, resulting in greater variability in responses within that personality category. Similarly, Salecha et al. (2024) show that models skew responses to socially desirable answers for the Big Five Personality test when they infer that they are evaluated. We demonstrate that social desirability bias also appears in other evaluation dimensions. This social desirability tendency explains the model’s difficulty to adhere to the unconscientious and antagonistic personas.
@@
The off-diagonal elements in the subtables of Table 1 reveal that most spill-over consistency effects occur across two evaluation categories: happiness and personality. For example, the off-diagonal entropy scores for Llama-3.3 70B are lower for the happiness and personality categories compared to occupation and political stance. This suggests that the model is more consistent across these two categories when they are not the assigned persona. Interestingly, this pattern differs from the intra-persona results, where happiness and occupation showed the strongest consistency when they were the assigned persona. The occupation and political categories are harder for the models to express, as not adding those personas results in responses without any occupational information or political stance. Manual analysis reveals how these personas are less frequently and explicitly expressed, especially in conversational settings. For instance, political beliefs rarely surface in responses to questions like "What are your music preferences?". Similarly, occupation-related information rarely appears unless explicitly prompted, though it may occasionally surface in essay-style or social media posts. The other two categories show spill-over effects. Additional manual analysis reveals that happiness and personality directly influence linguistic style: models default to a positive tone unless instructed otherwise, making happiness more overt. Similarly, personality traits, like extroversion, shape response style, amplifying spillover effects.
@@
Figure 2 shows how a sad persona is portrayed as more introverted, less conscientious, and less open than a happy persona. The economically right-winged and socially authoritarian personas are both less agreeable than their counterparts. All occupation personas are presented as extroverts, except the economist, who is more introverted. These observations reinforce prior findings that persona-assigned LLMs are susceptible to stereotypes Gupta et al. (2023). We show that these stereotypes also appear across general text-generation tasks. Furthermore, the model tends to answer in a happy, conscientious, agreeable, and open manner unless otherwise instructed or influenced by a stereotype. This is reflected in the heatmaps: the rows corresponding to the happiness, conscientiousness, and agreeableness evaluation categories generally lean toward a value of 1, indicating strong alignment. In contrast, neuroticism tends to lean toward 0, suggesting lower identification with that trait. This again illustrates the social desirability bias in the models  Salecha et al. (2024). Additionally, we show how personas can partially counteract this bias, e.g. the high consistency scores for neurotic and sad personas. However, this effect is not universal, as evidenced by the lower intra-persona consistency for antagonistic and unconscientious personas. Furthermore, most models tend to be slightly economically left and socially libertarian, though this varies by model and sometimes leans toward inconsistency. These default personas reveal the LLM’s ideological stances, shaped by training data and choices from model developers, called design choices Buyl et al. (2025); Cambo and Gergle (2022).
@@
Figure 3 shows that the survey, essay, and social media dimensions have the lowest intra- and inter-persona entropy, indicating the highest consistency across these tasks. However, consistency scores vary notably across dimensions, with mostly low correlations between them. This highlights the importance of considering all three dimensions to fully capture model behavior. Of these, only the inter-persona consistency between essay and social media is strongly correlated. Their low entropy scores could be attributed to the clarity of the task, where models can easily express their assigned persona. As tasks become less straightforward—such as answering open-ended questions about music preferences (singlechat and multichat)—models generally show a decrease in both intra- and inter-persona consistency. Moreover, the difference between intra- and inter-persona entropy becomes smaller. These results suggest that as tasks require less explicit persona expression, models may struggle to express distinct persona characteristics. Depending on the application, this variability may be advantageous, i.e. in creative or open-ended generation tasks where diversity is desirable. However, in more controlled settings that require reliable persona adherence, such unpredictability can be a drawback, as the model’s outputs may deviate from the intended persona or produce inconsistent behavior. Interestingly, the complexity of multichat scenarios compared to singlechat does not appear to hinder consistency. Contrarily, consistency increases slightly as follow-up responses allow models to provide more information, expressing the assigned persona more clearly. Nevertheless, consistency scores between singlechat and multichat are highly correlated. In addition to the main results, supplementary statistical analyses revealed no significant difference between inter- and intra-persona consistency in the singlechat and multichat dimensions. However, we observed statistically significant differences between the evaluation dimensions of survey, essay, and social media on the one hand, and singlechat (for both intra- and inter-persona consistency) as well as multichat (for intra-persona consistency) on the other. Concerning inter-persona consistency specifically, only the social media post generation task significantly outperforms the inter-persona consistency observed in the multichat evaluation dimension. Detailed statistical analysis is reported in Appendix I.